OpenStudy (anonymous):

What is the Integral for cos3xdx??

OpenStudy (amistre64):

seperate it into sin and cos parts ...

OpenStudy (amistre64):

im assuming that cos^3 (x) hard to read when people dont make use of grouping symbols

OpenStudy (anonymous):

Sorry, it's cos(3x)dx....

OpenStudy (amistre64):

ah, even simpler then

OpenStudy (amistre64):

we know that this had to come from some version of sin(3x) right?

OpenStudy (anonymous):

Okay

OpenStudy (amistre64):

lets tack on a constant scalar:\[k~\sin(3x)\] and take the derivative to compare it against what you have

OpenStudy (amistre64):

what would we get for this derivative?

OpenStudy (anonymous):

What about 3 sin (3x) ?

OpenStudy (amistre64):

hmm, lets try that out: 3 sin(3x) derives to 3*3 cos(3x) = 9 cos(3x) 9 cos(3x) is not equal to cos(3x)

OpenStudy (anonymous):

Mmmm, then 1/3 sin(3x) I presume?

OpenStudy (amistre64):

if we take the derivative of: k sin(3x) we obtain k sin(3x) derives to 3k cos(3x) 3k cos(3x) = cos(3x) when 3k = 1 so yeah :)

OpenStudy (anonymous):

Awesome!!! Thanks a lot amistre64!!!

OpenStudy (amistre64):

youre welcome